OpenStudy (anonymous):

determine the gragh of y=x^2-6x+3 has a maximun or minimum point then find the max or min point value. I think my is correct (-6,3) am I right

OpenStudy (anonymous):

\[y'=2x-6\] put y'=0 x=3 \[y''=2>0\]\[y(3)=9-18+3=-6\] so (3,-6) is minimum value.